OpenStudy (anonymous):

Choose the correct equation for the function of the graph shown below. y = 1/2cos1/4x y = 1/4sinx1/2 y = 1/2 cos1/4x y = 1/4sin1/2x

OpenStudy (anonymous):

heres the graph

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hello

OpenStudy (misty1212):

first lets decide sine or cosine any ideas? hint: \(\sin(0)=0,\cos(0)=1\)

OpenStudy (anonymous):

im really not sure

OpenStudy (misty1212):

take a look at the graph what is y when x is 0?

OpenStudy (anonymous):

i dont see when x is 0

OpenStudy (misty1212):

????

OpenStudy (misty1212):

where is 0 on the x axis?

OpenStudy (anonymous):

in the middle of the crossing lines

OpenStudy (misty1212):

yes at the origin \((0,0)\) so asking what is the curve at \(x=0\) is asking where does it cross the \(y\) axis

OpenStudy (anonymous):

i guess it'll on 0.5

OpenStudy (misty1212):

yes i think so too

OpenStudy (misty1212):

which tells you two things first it cant be sine, since \(\sin(0)=0\)

OpenStudy (misty1212):

so it must be cosine and also since \(\cos(0)=1\) and here you have \((0.\frac{1}{2})\) the amplitude is \(\frac{1}{2}\) so there must be a \(\frac{1}{2}\) out front

OpenStudy (misty1212):

so the answer is either A or C they look identical to me so maybe there is a typo somewhere in the question

OpenStudy (anonymous):

i didin't notice.... its the question because thats how its on my paper